1. How do higher organisms obtain ammonia?

2. Why is arginine classified as an essential amino acid in humans when it is synthesized in the urea cycle?

3. What is the fate of proteases after their role in protein digestion?

4. What would be the metabolic consequences of cells not expressing LDL receptors?

5. Vitamin D3 is derived from cholesterol. How and in which tissues does conversion occur?

6. How much energy, in the form of ATP, is obtained from b-oxidation of 1 mol of linoleic acid?

 


1
Expert's answer
2021-05-06T08:40:29-0400

1. Higher organism (complex organisms) acquire ammonia in tow ways; one is through nitrogen fixation, in which nitrifying organism such a nitrobacter residing in roots nodules of legumes e.g Rhizobium, converts nitrogen gas from the atmosphere in the soil to nitrates (ammonium nitrates). Another way in which ammonia is availed to higher organisms is through decomposition process; the decomposers dissolves tissues of dead plants and animals and releases the nutrients including ammonium nitrates into soil and eventually get absorbed by plants and other organisms such as fungi.


2. Arginine (2 amino-5-guanidinovaleric acid) is one of the 20 amino acid that are coded for a part of ribosomal protein synthesis in humans. When arginine is synthesized in the urea cycle, it not only act as a chemical intermediate in the release of nitrogenous waste, where the enzyme arginase (ARG) converts ornithine and urea allowing it (urea) to be secreted; arginine is also a structural component of many proteins and also serves as therapeutic target in a number of diseases, wound healing as well as cancer.


3. Once the digestion (hydrolysis) of proteins is complete, the residual enzyme (protease) are cleaved by other proteases, mostly of the same variety resulting in a regulation mechanism of protease regeneration, in order words, residual proteases are "recycled".


4. The primary role of LDL Receptors is to keep the amount of cholesterol in the blood at normal level. Malfunctioning or failure of the cell to express LDL receptors results in an increase or build up of cholesterol and accelerates atherosclerosis in human while in rabbits, it leads to hypercholesterolaemia and atherosclerosis.


5. The epidermis is the major source of vitamin D for the body. Under the influence of sunlight (ultraviolet radiation, action spectrum 280–320nM or UVB) 7-dehydrocholesterol in the epidermis is converted to vitamin D.


6. Number of FADH2 = (n/2) - 1 - 2 double bonds

= (18/2) -1 -2

= 9 -3

= 6

"\because" 8NADH + 6FADH2 + 9Acetyl CoA

No. of ATP = 8(2.5) + 6 (1.5) + 9(10)

= (20 + 9 + 90)ATP

= 119ATP